# A two-circuit approach to reducing quantum resources for the quantum lattice Boltzmann method

### v2

@@ Abstract
Computational fluid dynamics (CFD) simulations often entail a large computational burden on classical computers. At present, these simulations can require up to trillions of grid points and millions of time steps. To reduce costs, novel architectures like quantum computers may be intrinsically more efficient at the appropriate computation. Current quantum algorithms for solving CFD problems use a single quantum circuit and, in some cases, lattice-based methods. We introduce the a novel multiple circuits algorithm that makes use of a quantum lattice Boltzmann method (QLBM). The two-circuit algorithm we form solves the Navier–Stokes equations with a marked reduction in CNOT gates compared to existing QLBM circuits. The problem is cast as a stream function–vorticity formulation of the 2D Navier–Stokes equations and verified and tested on a 2D lid-driven cavity flow. We show that using separate circuits for the stream function and vorticity lead to a marked CNOT reduction: 35% in total CNOT count and 16% in combined gate depth. This strategy has the additional benefit of the circuits being able to run concurrently, further halving the seen gate depth. This work is intended as a step towards practical quantum circuits for solving differential equation-based problems of scientific interest.
@@
00footnotetext: Code available at: https://github.com/comp-physics/QLBM-frugal
I Introduction
Quantum computers can achieve complexity improvements by leveraging quantum mechanics principles such as superposition and entanglement. One example of quantum computing overcoming the limitations of computation on classical computers is Shor’s algorithm [1], which computes the prime factorization of integers in logarithmic time, providing an exponential speedup. Efforts are being made to explore the possibility of using quantum resources to speed up other important tasks. Example tasks include linear systems solvers [2, 3], Monte Carlo methods [4], and machine learning algorithms [5, 6]. With the recent progress of quantum hardware, various quantum algorithms have been purposed and examined for optimization problems [7], quantum chemistry [8], and finance [9].
An emerging use case of quantum computers is solving computational fluid dynamics (CFD) problems [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. For example, some techniques for solving the incompressible Navier–Stokes equations rely on an efficient on a sparse linear solver. The Harrow–Hassidim–Lloyd (HHL) algorithm [2] promises an exponential speedup on such tasks, assuming the state preparation process and information readout are efficient. Lapworth [26] proposes a hybrid CFD solver based on the SIMPLE (Semi-Implicit Method for Pressure Linked Equations) algorithm and HHL. However, the number of operations HHL requires exceeds current quantum hardware capability. Instead, proposals focus on using variational quantum algorithms (VQA) as near-term replacements for HHL to solve linear systems without any guaranteed complexity advantage, which can sometimes even begin to account for nonlinaer effects [27]. For eample, Demirdjian et al. [28] solve the 1D advection-diffusion equation with a Carleman linearization technique and the variational quantum linear solver [29].
Mesoscale methods for solving the Navier–Stokes equations operate above the molecular level but below the continuum scale. These methods provide indirect solutions by solving the Boltzmann equation, which characterizes the statistical behavior of a system of particles. Two methods that can solve the Boltzmann equation are lattice gas automata [30, 31] and the lattice Boltzmann method (LBM) [32, 33]. Traditional LBM algorithms are suited for classical algorithms, though some algorithms for quantum devices have been developed. For example, previous art developed quantum algorithms for the lattice gas [34, 35, 36]. Itani and Succi [37] demonstrate a Carleman linearization for the collision term and further work in Itani et al. [38] shows unitary evolution for both collision and streaming operators. Further works have pushed these methods forward [39, 40].
@@ I Introduction
Budinski [41] introduced a quantum lattice Boltzmann method (QLBM) for the advection–diffusion equation. This method simulates particles across a grid of lattice sites and extracts macroscopic quantities from mesoscopic simulations. Budinski [42] adapted the quantum circuit approximation to the Navier–Stokes equations. Using the stream function–vorticity formulation of the Navier–Stokes equations, the pressure term is removed, and the problem is reduced to an advection–diffusion equation and a Poisson equation. The strategy of these works involves a single quantum circuit for both equations.
@@
This work formulates a two-circuit 2D QLBM to solve the incompressible Navier-Stokes equations. The method is validated against the classical LBM solution to the 2D lid-driven cavity flow problem. The quantum resources required for this QLBM approach are compared to other approaches. Section II describes the classical LBM theory for simulating the advection–diffusion equation. Section III.1 shows a single-circuit quantum implementation of the LBM for solving the advection–diffusion equation. In section IV, a description of the stream function-vorticity formulation of the Navier–Stokes equations and work on a two-circuit quantum implementation for solving the 2D Navier–Stokes equations are provided. Section V presents verification against the classically-solved LBM simulation for the lid-driven cavity problem and quantum resource estimation. Section VI serves to clarify the contributions of the method presented in this manuscript and its limitations.
II Classical lattice Boltzmann method
The LBM uses a grid of cells (shown in fig. 1) that contain particle distributions. The particles move into neighbor cells during each time step according to the Boltzmann equation [43, 44]:
@@ II Classical lattice Boltzmann method
We apply a new formulation to two PDE problems, the advection–diffusion and incompressible Navier–Stokes equations. Using the stream function–vorticity formulation, the Navier–Stokes equations reduce into two equations that closely resemble the advection–diffusion equation, so the latter will be examined first.
The advection–diffusion is
which represents advection as $c({\partial\phi}/{\partial x})$ and diffusion as $D({\partial^{2}\phi}/{\partial x^{2}})$. Here, $c$ and $D$ are advection and diffusion coefficients and $\phi(t,x)$ is a scalar concentration field.
For the advection–diffusion equation of (2), the equilibrium distribution function is
where $\phi(\bm{r},t)$ corresponds to the lattice site at lattice position $\bm{r}$ = $(x_{i},y_{j})$, $\bm{c}$ is the advection velocity vector, and $c_{s}=1/\sqrt{3}$ is the speed of sound for the D2Q5 scheme [46]. The notation $x_{i}$ and $y_{j}$ denotes the $i$-th cell in the $x$ direction and the $j$-th cell in the $y$ coordinate direction.
@@
Herein we use $D=1/6$, $c_{s}=1/\sqrt{3}$, and $\tau=\Delta t$. This gives $\varepsilon=\tau/\Delta t=1$ and
@@
which includes the collision (3) and the streaming steps (5). The concentration field is thus
@@
III Quantum lattice Boltzmann method
III.1 Advection–diffusion equation
A quantum circuit can implement the lattice Boltzmann method, which involves a similar approach to classical computation. Specifically, the quantum circuit performs the collision operator followed by the streaming of particles and subsequent recalculation of macroscopic variables. Boundary conditions are applied at the end of each time step. No special care is needed when boundary conditions are periodic; the left and right shift gates $R$ and $L$ automatically propagate boundary conditions to the designated lattice site for each link $\alpha$.
It is important to note that encoding input is a bottleneck in the algorithm. For the current study, the amplitude encoding technique described in the work of Shende et al. [47] is employed. Amplitude encoding is part of the Qiskit toolkit [48], though it is also a generic encoding algorithm with an expensive gate count.
We organize the quantum circuit for the advection–diffusion equation using a D2Q5 scheme with qubits organized into 4 registers, which are groupings of qubits. Registers $r_{0}$ and $r_{1}$ contain $\log M$ qubits, where $M$ is the number of lattice sites in each dimension. Register $d$ has $\lceil\log_{2}Q\rceil$ qubits, where $Q$ is the number of LBM links $\alpha$. Register $a$ holds a single ancilla qubit necessary for applying a non-unitary collision operator.
III.1.1 Encoding input
At the start of each time step, given a distribution $\phi(\bm{r},t)$, define $\phi(\alpha,\bm{r})$ to be the value of $\phi$ at time $t$ and position $\bm{r}$ at link direction $\alpha$. Given a discretized concentration
@@ III.1.1 Encoding input
This normalizes $\phi(\alpha,\bm{r})$, and the initial qubit states follow from amplitude encoding. In fig. 2, register $r_{0}$ stores the data at each link direction $\alpha$, and it can be retrieved by specifying the link $\alpha$ in register $d$.
@@
which follow from (3). This strategy is discussed further in the next subsection.
III.1.2 Collision operator
Figure 2: QLBM circuit for an advection–diffusion time step. The collision operator $A=(C_{1}+C_{2})/2$, and shift operators $R$, $L$, $D$, and $U$ propagate particles in each link direction $\alpha$. In panel (c), a Hadamard [49] gate is applied to the qubits in register $d$.
@@ III.1.2 Collision operator
The coefficient matrix $A$ is not unitary, so it cannot be directly translated into a quantum gate. An important strategy was described by Budinski [42]: A linear combination of unitary matrices split $A$ into two matrices $C_{1}$ and $C_{2}$, related to the original matrix as
@@
The circuit of fig. 3 evolves an input statevector |$\ket{\psi}$ according to the unitary
@@
are the link coefficients described by (10). Panel (a) of fig. 2 uses a linear combination of unitaries to apply matrix $A$ to the input. The Hadamard gates are used for the block encoding process along with $C_{1}$ and $C_{2}$ operations, which are derived from the unitary matrices of (11). The coefficients representing the proportion of particles in each link $\alpha$ are $k_{1}$ and $k_{2}$. The collision matrices are
$C_{1,2}=\begin{bmatrix}\exp(\pm i\cos^{-1}(k_{1}))I_{n}&0\\ 0&\exp(\pm i\cos^{-1}(k_{2}))I_{n}\end{bmatrix}.$ (20)

### v3

# A multiple-circuit approach to quantum resource reduction with application to the quantum lattice Boltzmann method

@@ Abstract
This work proposes a multi-circuit quantum lattice Boltzmann method (QLBM) algorithm that leverages parallel quantum computing to reduce quantum resource requirements. Computational fluid dynamics (CFD) simulations often entail a large computational burden on classical computers. At present, these simulations can require up to trillions of grid points and millions of time steps. To reduce costs, novel architectures like quantum computers may be intrinsically more efficient for these computations. Current quantum algorithms for solving CFD problems are based on single quantum circuits and, in many cases, use lattice-based methods. Current quantum devices are adorned with sufficient noise to make large and deep circuits untenable. We introduce a multiple-circuit algorithm for a quantum lattice Boltzmann method (QLBM) solve of the incompressible Navier–Stokes equations. The method, called QLBM-frugal, aims to create more practical quantum circuits and strategies for differential equation-based problems. The presented method is validated and demonstrated for 2D lid-driven cavity flow. The two-circuit algorithm shows a marked reduction in CNOT gates, which consume the majority of the runtime on quantum devices. Compared to the baseline QLBM technique, a two-circuit strategy shows increasingly large improvements in gate counts as the qubit size, or problem size, increases. For $64$ lattice sites, the CNOT count was reduced by 35%, and the gate depth decreased by 16%. This strategy also enables concurrent circuit execution, further halving the seen gate depth.
@@
Code available at: https://github.com/comp-physics/QLBM-frugal
1 Introduction
1.1 Motivation and application of quantum computation for scientific computing
Computationally intensive algorithms are obstacles to the scientific computing and engineering communities. Quantum computers can achieve complexity improvements by leveraging quantum mechanics properties, making them of interest. A prominent example of this so-called quantum advantage is demonstrated by Shor’s algorithm [1]. The algorithm finds the prime factorization of integers in logarithmic time, an exponential speedup over the optimal classical algorithm. Other quantum algorithms include linear systems solvers [2, 3], Monte Carlo methods [4], and machine learning algorithms [5, 6]. With the recent progress of quantum hardware, these algorithms have been applied in industries involving optimization [7], quantum chemistry [8], and finance [9]. The potential for such large speedups makes partial differential equation (PDE) solvers attractive. However, classical problems in this realm tend to be nonlinear and non-unitary, presenting hurdles for quantum computing [10]. Attention has been turned toward addressing this challenge.
1.2 Quantum computation for computational fluid dynamics
Computational fluid dynamics (CFD) problems often require some of the world’s largest classical supercomputers to solve. Focus has been directed toward alternative quantum algorithms for these problems [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. The connection between quantum mechanics and fluid dynamics emerged early on through the Madelung equations, a reformulation of the Schrödinger equations [27, 12] This relationship between quantum computing and CFD was further explored in 1993, when Succi and Benzi [28] used the lattice Boltzmann equation to describe non-relativistic quantum mechanics.
Over a decade later, the Harrow–Hassidim–Lloyd (HHL) algorithm, a sparse linear solver with quantum advantage, was used to solve the incompressible Navier–Stokes equations [2]. Many quantum CFD solvers are also based on the quantum singular value transformation algorithm [29]. Hybrid quantum and classical algorithms have also been presented for CFD problems [30, 31, 32]. One reduces the CFD problem to a quantum-solvable form, and the quantum device solves the linear system. However, the advantage posed by the HHL algorithm does not account for the state preparation and readout requirements [33]. Implementing efficient input and output methods for quantum computers is in itself a challenge: loading an exponentially large state space representative of $n$-qubits requires $2^{n}$ amplitudes be somehow encoded on a quantum computer. Various approaches have been taken, but present generalized encoding methods are upper-bounded by an exponential complexity [34, 35, 36], with slight improved efficiency and depth on encoding sparse matrices [37].
1.3 Quantum computing for mesoscale CFD methods
This paper extends bodies of work focused on mesoscale methods for fluid simulations. The dissipative particle dynamics and Boltzmann equation-based methods are commonly used in classical computation. The Boltzmann equation (which characterizes the statistical behavior of a system of particles) is focused on herein. This equation is primarily solved using lattice gas automata [38, 39] and the lattice Boltzmann method (LBM) [40, 41]. LBM is the more advantageous of the two, because it enables noise resilience and flexibility in complex multiphysics simulations. Prior bodies of work have presented adaptations of both methods. For example, previous bodies of work proposed quantum algorithms for the lattice gas model [42, 43, 44]. More recently, a quantum lattice gas model was presented by Zamora, Budinski, Niemimaki, and Lahtinen [45] with improved efficiency. In 2020, Todorova and Steijl [46] described a quantum algorithm for the collisionless Boltzmann equation. This method accounts for the case of nearly free molecular flows and ignores complex particle collisions.
Linearizing the problem is the first step in mapping the lattice Boltzmann method to a quantum-suitable form. This struggle with nonlinear terms is due to the linear framework on which quantum computing is defined. The Carleman and Koopman–von Neumann embeddings have been used to linearize equations for quantum computing [47]. The Carleman method embeds the variables of nonlinear equations directly into a quantum state. This approach is advantageous over the Koopman–von Neumann method, which instead embeds the variables through orthogonal polynomials. Itani and Succi [48] demonstrated a Carleman linearization for the collision term and later extended the work to demonstrate unitary evolution for both the collision and streaming operators [10]. These approaches were extended in later work [49, 46].
@@ 1.3 Quantum computing for mesoscale CFD methods
Budinski [50] employed the Carleman linearization technique to introduce a quantum lattice Boltzmann method (QLBM) for the advection–diffusion equation. The proposed method simulates particles across a grid of lattice sites and extracts macroscopic quantities from mesoscopic simulations. Since the resultant matrix is not necessarily Hermitian, the system must be discretized in terms of time [2, 51]. The linearized system of equations is then solved via a quantum linear system algorithm [47]. Sanavio and Succi [49] introduced a QLBM algorithm that simulated Kolmogorov-like flows using the Carleman linearization method. They showed that their implementation of the single time-step operator would have a fixed circuit depth. Budinski [52] adapted the quantum circuit approximation using the stream function–vorticity formulation of the Navier–Stokes equations. This adaptation removes the pressure term, reducing the problem to the advection–diffusion equation and Poisson equation, solved on a single circuit.
1.4 Limitations of quantum computing
Many of the proposed algorithms are presumed to operate on fault-tolerant quantum computers [52, 49, 17, 53]. Although the theory is promising, noise and decoherence on present quantum devices result in high error rates. Consequently, foundational algorithms, such as HHL, are unable to be supported on modern quantum hardware [30]. The coherence time of a device is the range of time during which the quantum device reliably maintains its state such that the results obtained are within reason [54, 55]. Circuits whose runtime exceeds this coherence time output solutions that contain too much noise to be useful. The sensitivity to the environment makes it difficult to scale quantum algorithms, as they require large numbers of qubits. As a result, the aforementioned quantum CFD algorithms must be simulated.
The limitations in current quantum hardware is denoted Noisy Intermediate-Scale Quantum (NISQ). These limitations have driven research towards hybrid variational quantum algorithms (VQAs) for both adiabatic and gate-based quantum computing [56]. The adiabatic configuration models the system as an energy minimization problem cast as an Ising Hamiltonian [57]. However, this approach is unable to adapt to allow control over precision and other parameters [58].
Gate-based VQAs are proposed as near-term replacements for HHL without the quantum advantage. Demirdjian et al. [59] solved the 1D advection–diffusion equation via Carleman linearization with reasonable accuracy. This approach uses a variational quantum linear solver [60] to compute the solution. Lubasch et al. [61] demonstrated VQAs may account for nonlinear effects. However, beyond a certain circuit depth in noisy devices, VQAs are unlikely to outperform optimal classical counterparts on combinatorial problems [62]. Thus, methods to reduce the circuit depth and number of quantum gates applied are beneficial.
1.5 Distributed and parallelized quantum systems
One approach to reducing quantum resource use is distributed computation. A distributed quantum computer is a novel architecture composed of a network of devices. Its interactions occur via quantum teleportation, physical transport, or a predefined classical method. Attention has been dedicated toward nonlocal quantum operations that enable their communication [63, 64, 65]. Distributed computing is of interest in quantum communications, quantum internet, and quantum key distribution research, among others [66, 67, 68]. This architecture has been conceptualized as a technique to adapt to limitations on small capacity quantum computers, including on the foundational Simon’s and Grover’s algorithms [63, 69, 70]. Sub-processes within the algorithms are delegated among the devices.
There has also been interest in parallelized quantum processes. In this architecture, multiple circuits are run concurrently on a single device. Broadbent and Kashefi [71] devises a technique for parallelizing quantum circuits that reduces polynomial-depth circuits to logarithmic depth. The quantum processor’s qubits are allocated to minimize interference from the concurrent processes. Matteo and Mosca [72] demonstrated that this method reduces gate depth and speedup in select cases. As such, investigating the parallelization of proposed quantum algorithms is beneficial when solving larger problems.
1.6 Objectives
@@
This work formulates a two-circuit 2D QLBM algorithm, QLBM-frugal, to solve the incompressible Navier-Stokes equations. The method is validated against the classical LBM solution to the 2D lid-driven cavity flow problem. The quantum resources required for QLBM-frugal are compared to other approaches, and the parallelized implementation is discussed. Section 2 describes the classical LBM theory for simulating the advection–diffusion equation. Section 3 shows a single-circuit quantum implementation of the LBM for solving the advection–diffusion equation. In section 4, the two-circuit quantum implementation for solving the 2D Navier–Stokes equations is provided. Section 5 presents verification against the classically-solved LBM simulation and quantum resource estimations. Section 6 discusses the current limitations of the work. Section 7 summarizes the contributions of the presented method and its potential impact.
2 Mathematical formulation
2.1 Advection–diffusion equation
The advection–diffusion equation describes flow in a way that treats advection and diffusion as concurrent processes. The governing advection–diffusion PDE is
where advection is defined by $c({\partial\phi}/{\partial x})$ and diffusion is defined by $D({\partial^{2}\phi}/{\partial x^{2}})$. Here, $c$ and $D$ are the advection and diffusion coefficients, and $\phi(t,x)$ is a scalar concentration field. We use the advection–diffusion algorithm devised by Budinski [50] as the foundation for the presented multi-circuit approach.
2.2 General lattice Boltzmann formulation
The LBM model for fluid flow simulates the evolution of particle flow over time, given by [73, 74] as
@@ 2.2 General lattice Boltzmann formulation
where $f_{\alpha}$ indicates the particle distribution along each $\alpha$. $e_{\alpha}$ is the velocity of a particle in link $\alpha$, $t$ is time, $\delta t$ is the time step, $\bm{r}$ is the cell position, $S$ is the source term, $w_{\alpha}$ is the proportion of particles streaming in link $\alpha$, and $\epsilon=\Delta t/\tau$ where $\tau$ is the relaxation time. We more generally define $\bm{r}$ using Cartesian coordinates, whose notation depends on the dimension of the vector space. For example, in a one-dimensional space, we express $\bm{r}=(x)$, while in a two-dimensional space, we express $\bm{r}=(x,y)$. Figure 1 shows how the particle distribution may be visualized along a grid of cells. Each time step indicates some particle movement into neighboring cells in accordance with the above equation.
@@
2.3 Formulation for one-dimensional scheme
Initially, we devise a one-dimensional model for the D1111Q2222 configuration. Two velocity vectors are considered such that
$e_{1}=-e_{2}=\Delta x/\Delta t$ (3)
for corresponding distribution functions $f_{1}$ and $f_{2}$. The symmetric boundary condition of the LBM imposes this equality. Our model resolves around the local equilibrium distribution function [75], formulated as
where $\phi(\bm{r},t)$ corresponds to the lattice site at lattice position $\bm{r}$ = $(x_{i})$, $\bm{c}$ is the advection velocity vector, and $c_{s}$ is the speed of sound. We set the weights $w_{1,2}=0.5$ and $c_{s}=1$, in accordance with the approach by Budinski [50]. Likewise, the D1111Q3333 configuration is considered with three velocity vectors, where $e_{0}=0$ and $e_{1}=-e_{2}=1$. Let the speed of sound $c_{s}=1/\sqrt{3}$ following Servan-Camas and Tsai [76].
2.4 Formulation for two-dimensional scheme
We now consider the two-dimensional D2222Q5555 scheme. Budinski [52] shows that these equations are reduced to resemble the advection–diffusion formulation, referencing his earlier work [50] in setting the parameters and governing formulas. We adapt this advection–diffusion formulation to a multiple-circuit method for advection and diffusion. This formulation is expressed as
where advection is defined by $c({\partial\phi}/{\partial x})$ and diffusion is defined by $D({\partial^{2}\phi}/{\partial x^{2}})$. Here, $c$ and $D$ are the advection and diffusion coefficients and $\phi(t,x)$ is a scalar concentration field.
The equilibrium distribution function is solved using 4, where $\bm{r}=(x_{i},y_{i})$. Let $c_{s}=1/\sqrt{3}$ [76] and the weights be defined such that $w_{0}=2/6$ and $w_{2,3,4}=1/6$ respective to the directions illustrated in fig. 1 [77].
@@ 2.4 Formulation for two-dimensional scheme
Herein we use $D=1/6$ and $\tau=\Delta t$. Recall we have set $c_{s}=1/\sqrt{3}$. This gives $\varepsilon=\tau/\Delta t=1$ and
@@
which includes the collision 4 and the streaming steps 7. The concentration field is thus
@@
3 Quantum Lattice Boltzmann method
3.1 Quantum Lattice Boltzmann method for the advection–diffusion equation
We now discuss the premise of the body of work devised by Budinski [52]. Consider, first, the advection–diffusion equations. The quantum circuit performs the collision operator, followed by the streaming of particles and recalculation of macroscopic variables. Boundary conditions are applied at the end of each time step. No special care is needed when boundary conditions are periodic; the left and right shift gates $L$ and $R$ automatically propagate boundary conditions to the designated lattice site for each link $\alpha$. Figure 2 illustrates how the quantum circuit’s qubits are organized into 4 registers, or groupings of qubits, to solve the advection–diffusion equation using a D2Q5 scheme. Registers $r_{0}$ and $r_{1}$ contain $\log N_{\mathrm{site.}}$ qubits, where $N_{\mathrm{site.}}$ is the number of lattice sites in each dimension. Register $d$ has $\lceil\log_{2}Q\rceil$ qubits, where $Q$ is the number of LBM links $\alpha$. Register $a$ holds a single ancilla qubit necessary for applying a non-unitary collision operator.
3.1.1 Encoding input
For the current study, the amplitude encoding technique described in the work of Shende et al. [78] is employed. Amplitude encoding is part of the Qiskit toolkit [79], albeit also a generic encoding algorithm with an expensive gate count. We further elaborate on this in section 6.
At the start of each time step, we are given a distribution $\phi(\bm{r},t)$ We define $\phi(\alpha,\bm{r})$ to be the value of $\phi$ at time $t$ and position $\bm{r}$ at link direction $\alpha$. Given a discretized concentration
@@ 3.1.1 Encoding input
This formulation normalizes $\phi(\alpha,\bm{r})$, and the initial qubit states follow from amplitude encoding. In fig. 2, register $r_{0}$ stores the data at each link direction $\alpha$. This data can be retrieved by specifying the link $\alpha$ in register $d$.
@@
which follow from 4. This strategy is discussed further in the next subsection.
3.1.2 Collision operator
Figure 2: QLBM circuit for an advection–diffusion time step as devised by Budinski [50]. The collision operator $A=(C_{1}+C_{2})/2$, and shift operators $R$, $L$, $D$, and $U$ propagate particles in each link direction $\alpha$. In panel (c), a Hadamard [80] gate is applied to the qubits in register $d$.
@@ 3.1.2 Collision operator
The coefficient matrix $A$ is not unitary, so it cannot be directly translated into a quantum gate. To mitigate this, an important strategy, devised by Childs and Wiebe [81], is employed by Budinski [52]. Suppose we split $A$ into a linear combination of unitary matrices, $C_{1}$ and $C_{2}$, related to the original matrix as
@@
The circuit in fig. 3 evolves an input statevector |$\ket{\psi}$ according to the unitary
@@
are the link coefficients described by 12. Panel (a) of fig. 2 uses a linear combination of unitaries to apply matrix $A$ to the input. The Hadamard gates are used for the block encoding process along with $C_{1}$ and $C_{2}$ operations, which are derived from the unitary matrices of 13. The coefficients representing the proportion of particles in each link $\alpha$ are $k_{1}$ and $k_{2}$. The collision matrices are
$C_{1,2}=\begin{bmatrix}\exp(\pm i\arccos{(k_{1})})I_{n}&0\\ 0&\exp(\pm i\arccos{(k_{2})})I_{n}\end{bmatrix}.$ (16)